OpenStudy (anonymous):

Can someone clarify what happens to the negative sign when we have -(sqrt3/2)^2 + 1 I know it equals 1/2, but i'd like to know what happens to the negative sign. Thx

OpenStudy (anonymous):

\[(-\sqrt{3/2} )^{2}+1\] ,is that the equation ????

OpenStudy (anonymous):

yup Eyad

OpenStudy (anonymous):

then cause of the power if the power is even so it delete the -ve sign ,while if its is odd i dont delete the power

OpenStudy (anonymous):

\[-\sqrt{\frac{3}{2}}^2+1=-\frac{3}{2}+1=-\frac{1}{2}\] whereas \[\left(-\sqrt{\frac{3}{2}}\right)^2+1=\frac{3}{2}+1=\frac{5}{2}\]

OpenStudy (anonymous):

if u need more explanation ,tell me :)

OpenStudy (anonymous):

Maybe i should of put the whole thing down, because it says: −(3√2)2+1=cos2x, cos x = 1/2. Sorry

OpenStudy (anonymous):

I got -1/2 too, but i'd like to know what happens to make it positive?? thx

OpenStudy (anonymous):

\[−(√3/2)^2+1=\cos2x\] cos 2x = 1/2

OpenStudy (anonymous):

\[−(√3/2)^2+1=\cos2x\] cos x = 1/2